OpenStudy (anonymous):

help!!!!! (a+b)^3=??? (a+b)^4=??? (a+b)^5=???

OpenStudy (kainui):

What does the ^3 or ^4 part really mean? Surely if you see x^3 you can rewrite it as x*x*x right?

OpenStudy (anonymous):

^3 means to the third power

OpenStudy (anonymous):

yeah

OpenStudy (kainui):

So rewrite it like that without exponents and see if you can do it now.

jhonyy9 (jhonyy9):

(a+b)^3 = (a+b)(a+)^2 =(a+b)(a^2 +2ab +b^2) = hope you can continue it --- so the rest two you can calculi the same too hope so much that you will understand it all right good luck bye